OpenStudy (anonymous):

Explain, in complete sentences, how to use the distributive property to simplify the expression below. 3(4 + x)

jimthompson5910 (jim_thompson5910):

take the outer number 3 and multiply it by EVERY number or term inside so.... 3 times 4 = 12 3 times x = 3x

jimthompson5910 (jim_thompson5910):

now add these products: 12 + 3x

jimthompson5910 (jim_thompson5910):

this means 3(4 + x) turns into 12 + 3x

OpenStudy (anonymous):

gee thx can u help me with more

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

thx wait a min

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

this is a fraction -2x - 8/ -2

jimthompson5910 (jim_thompson5910):

so it's \[\Large \frac{-2x-8}{-2}\] right?

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

ok so you use the distributive property...but in reverse

OpenStudy (anonymous):

ok so is it x + 4

jimthompson5910 (jim_thompson5910):

think of -2x - 8 how can you factor that?

jimthompson5910 (jim_thompson5910):

it is x+4, but it helps to know why it's x+4

OpenStudy (anonymous):

im lost, do you multiply it

OpenStudy (anonymous):

at firts i got x- 8

jimthompson5910 (jim_thompson5910):

so sorry, i lost connection

jimthompson5910 (jim_thompson5910):

-2x - 8 factors to -2(x+4) the -2 terms cancel, so that's why the answer is x+4

jimthompson5910 (jim_thompson5910):

do you see how to get this?

OpenStudy (anonymous):

thxs @jim_thompson5910 i need more help :) Use the distributive property to simplify the expression below: 5t − 7(3 + t)

jimthompson5910 (jim_thompson5910):

first distribute − 7(3 + t)

jimthompson5910 (jim_thompson5910):

so -7 times 3 = -21 -7 times t = -7t

jimthompson5910 (jim_thompson5910):

which means 5t − 7(3 + t) turns into 5t − 21 - 7t

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

dont u have to do add 5t and 71 since they are both like terms or

jimthompson5910 (jim_thompson5910):

you mean 5t and -7t right?

OpenStudy (anonymous):

or do you subtract it

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

yes 5t - 7t = ???

OpenStudy (anonymous):

-2t

OpenStudy (anonymous):

so is the answer -2t - 21

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yup i get one more help plz

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

3x-21/ 3

OpenStudy (anonymous):

its fraction

jimthompson5910 (jim_thompson5910):

ok use distribution in reverse to go from 3x - 21 to 3(x-7)

jimthompson5910 (jim_thompson5910):

notice how if you distribute again, you'll get 3x - 21 back

OpenStudy (anonymous):

oh at first i got x -21

OpenStudy (anonymous):

or is it 3x - 7

jimthompson5910 (jim_thompson5910):

no it's neither of those

jimthompson5910 (jim_thompson5910):

the 3s will cancel leaving you with x-7

OpenStudy (anonymous):

ohhhhh, i get it you cancel out all the 3 then the 7 takes the negative sign oof 21

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

one more its pretty easy, i know the algebretic terms, but this is confusing The difference of 3 times 4 and 2

OpenStudy (anonymous):

what does it equal

jimthompson5910 (jim_thompson5910):

difference means subtraction, so The difference of 3 times 4 and 2 means 3*4 - 2

OpenStudy (anonymous):

lol, i got 10 at first, im this type of paranoid person that i have to get everything right in my homework, so i get it corrected in openstudy, im in 4th grade advanced class ge thx for all ur help

jimthompson5910 (jim_thompson5910):

well 3*4 = 12 so 3*4 - 2 is the same as 12 - 2

jimthompson5910 (jim_thompson5910):

and that's 10

jimthompson5910 (jim_thompson5910):

so if you wanted to simplify as much as possible, then it would be 10

OpenStudy (anonymous):

yup thx, i would really give u more medals if i can

jimthompson5910 (jim_thompson5910):

its fine, glad to be of help